OpenStudy (anonymous):

Show how to find the particular solution of the differential equation, dy/dx+5y=9, x=0 when y=1

OpenStudy (anonymous):

I could use a LaPlace transform right?

OpenStudy (anonymous):

Yeah.. and it's pretty easy.

OpenStudy (anonymous):

Ok the only problem I'm having is, all I have to use is the formula L(y')=sL(y)-y(0) correct?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

ok so I'm having trouble figuring out how to incorperate x=0 and y=1

OpenStudy (anonymous):

x=0 when y=1 means that y(0)=1

OpenStudy (anonymous):

Oh.. duh.

OpenStudy (anonymous):

so I would just rewrite... \[s L(y)-y(0)+5L(y)=L(9)\]

OpenStudy (anonymous):

OK.. what are the L(y) and L(9)?

OpenStudy (anonymous):

Not 100% sure on L(y) but I know L(9)=9/s

OpenStudy (anonymous):

Would it just be y/s ?

OpenStudy (anonymous):

or no, I could solve for L(y)

OpenStudy (anonymous):

I think I have it

OpenStudy (anonymous):

Yeah.. you could do so, or you could write it as Y(s)

OpenStudy (anonymous):

You know that you have to transform it back to the time domain.

OpenStudy (anonymous):

L(y)=(9/s+1)/(s+5)

OpenStudy (anonymous):

are you sure about the simplification you made?

OpenStudy (anonymous):

oh I see it.. it's right.

OpenStudy (anonymous):

I don't know how to do fractions notation using the equation thing or I would. :)

OpenStudy (anonymous):

I think I need to grab a pen and a paper!

OpenStudy (anonymous):

Well you so far have: \[Y(s)={9/s+1 \over s+5}\] you can multiply by s (since it's not 0).. you will get: \[{9+s \over s(s+5)}\]

OpenStudy (anonymous):

Oh right to clean it up. Ok so now how do I evaluate for the particular solution?

OpenStudy (anonymous):

\[{9+s \over s(s+5)}= {A \over s}+ {B \over s+5}\] how are you with partial fractions, can you find A and B?

OpenStudy (anonymous):

not good at ALL with partial fractions...

OpenStudy (anonymous):

it's easy.. multiply both sides of the last equation by s(s+5)

OpenStudy (anonymous):

You mean A(s(s+5)/s+B(s(s+5)/(s+5) ?

OpenStudy (anonymous):

Almost there!! something should be cancelled, right? And din't forget the left side of the equality sign.

OpenStudy (anonymous):

don't*

OpenStudy (anonymous):

A(s+5)+B(s)

OpenStudy (anonymous):

err

OpenStudy (anonymous):

that should equal to?

OpenStudy (anonymous):

9+s/s(s+5)=A(s+5)+B(s)

OpenStudy (anonymous):

nearly

OpenStudy (anonymous):

you multiply both sides by s(s+5)

OpenStudy (anonymous):

Oh right ok so its

OpenStudy (anonymous):

9+s=A(s+5)+B(s)

OpenStudy (anonymous):

Hey lokisan :).. sorry I took over. we seem to be useless wjen you're around :P

OpenStudy (anonymous):

you're doing fine :D

OpenStudy (anonymous):

I'm 'half' watching...

OpenStudy (anonymous):

there is more than one way find A and B now.. take s to be zero and find the value of A.

OpenStudy (anonymous):

If I did something wrong, Please let me know :)

OpenStudy (anonymous):

you're on the right track

OpenStudy (anonymous):

A=9/5

OpenStudy (anonymous):

Then I put that back into the original for B?

OpenStudy (anonymous):

or A and B for s...

OpenStudy (anonymous):

Yeah.. or just plug s=-5

OpenStudy (anonymous):

b=-4/5

OpenStudy (anonymous):

now what does Y(s) equal to?

OpenStudy (anonymous):

I'm kind of confused.. so I now put int he values of A and B back into 9+s=A(s+5)+B(s) to solve for the s?

OpenStudy (anonymous):

Do you know why we used the partial fractions in the first place?

OpenStudy (anonymous):

So we could evaluate what s was when x=0 and y=1

OpenStudy (anonymous):

No.. because we want to write an expression for Y(s) that can be easily transformed back to our original function.

OpenStudy (anonymous):

don't forget we are looking for the particular solution of the DE.

OpenStudy (anonymous):

My instructor never used the notation as Y(s) so I think thats where I'm losing you.

OpenStudy (anonymous):

Oh oh I think I understand... we're looking to change 9+s/s(s+5)

OpenStudy (anonymous):

\[L(y)={9+5 \over s(s+5)}={9 \over 5}.{1 \over s}-{4 \over 5}.{1 \over s+5}\]

OpenStudy (anonymous):

Isn't it 9+s

OpenStudy (anonymous):

oh yeah 9+s sorry

OpenStudy (anonymous):

now focus on the right part, and try to transform it back to y(x).

OpenStudy (anonymous):

Ugh I understand what you mean, I just am unaware of the method. I don't do inverse laplace right?

OpenStudy (anonymous):

You do nothing but the inverse Laplace.

OpenStudy (anonymous):

So like... 1/s=1 etc?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

not 1/s=1.. but the inverse laplace of 1/s=1 or more accurately u(x), (which is the unit step function)

OpenStudy (anonymous):

How do I solve the 9/5 and -4/5

OpenStudy (anonymous):

laplace transform and its inverse are both linear operations, that's L[af(t)]=aL[f(t)]

OpenStudy (anonymous):

Ugh I was understanding it and now I'm lost. I understand to use the table but this is just beyond me now.

OpenStudy (anonymous):

You should not worry about these constants.. just keep them as they are.

OpenStudy (anonymous):

Am I complicating things here? :(

OpenStudy (anonymous):

So just worry about the 1/s and 1/s+5?

OpenStudy (anonymous):

No you're doing good. I'm sure its just me.

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

you already did 1/s.. what about 1/(s+5)?

OpenStudy (anonymous):

I just think you're not familiar with notation.

OpenStudy (anonymous):

e^-at is the transform...

OpenStudy (anonymous):

e^-5t then right?

OpenStudy (anonymous):

You know what you need to do, though.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Me?

OpenStudy (anonymous):

No, Scotty.

OpenStudy (anonymous):

Exactly Scotty.

OpenStudy (anonymous):

Sorry...

OpenStudy (anonymous):

It's OK :)

OpenStudy (anonymous):

Ok so to clean it up I use use what I have left right?

OpenStudy (anonymous):

Which basically will end up e^5t

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Your answer should have e^(-5t) in it.

OpenStudy (anonymous):

9/5*1-4/5*e^(-5t)

OpenStudy (anonymous):

It's in x not in t.

OpenStudy (anonymous):

Yes, Scotty.

OpenStudy (anonymous):

But in x.

OpenStudy (anonymous):

I'm used to doing them in t...sorry.

OpenStudy (anonymous):

so e^-(5x)

OpenStudy (anonymous):

so final answer is: \[y(x)={9 \over 5}- {4 \over 5} e ^{-5x}\] right?

OpenStudy (anonymous):

We all are used to t.

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

Ok I was rushing through on the simplification.

OpenStudy (anonymous):

I'm 12 years old and what is this? \[\frac{\mathbb{d}y}{\mathbb{d}x} = 9 - 5y \implies \int \frac{1}{9-5y} \mathbb{d}y = \int \mathbb{d}x \implies c -\frac{1}{5} \ln 9 - 5y = x\] x = 0 when y = 1 \[\implies c = \frac{1}{5} \ln 4 \implies x = \ln \frac{4}{9-5y} \implies y = \frac{9-4e^{-5x}}{5}\]